1.Calculate the mass of oxide that can be obtained from 25.0 g of the metal.    

2.Calculate the number of moles of calcium chloride produced if 148 g calcium hydroxide is neutralized by hydrochloric acid.    Ca(OH)2(aq)+2HCL(aq)=CaCl2(aq)+2H2O(l)

           


1
Expert's answer
2021-04-25T03:50:15-0400

1) Let n=1 mol, then M(Metal)=25g/1mol=25g

The closet to this molar mass is Mg.

n(Mg)=m/M=25g/24g/mol=1.042 mol

2Mg+O2=2MgO

n(MgO)=n(Mg)=1.042 mol (by the equation)

m(Mg)=1.042mol*40g/mol=41.68g


2) M(Ca(OH)2)=40+1*2+16*2=74 g/mol

n(Ca(OH)2)=m/M=148g/74g/mol=2mol

n(CaCl2)=n(Ca(OH)2)=2 mol (by the equation)